You look through a Polaroid sheet at a 100 W light bulb 2.97 m away. (a) Find the intensity of the light you see. (b) Next you place a second Polaroid sheet in front of the first sheet with an angle of 44.1° between the two transmission axes. What intensity do you observe?
Answer #1

a)

intensity

I = P / (4 pi r^2)

I = 100 % ( 4* 3.14 * 2.97^2)

I = 0.9026 W/m^2

=====

b)

intensity after 1st polariser

I1 = I/2

intensity after 2nd polariser

I2 = (I/2) (cos 44.1)^2

I2 = 0.5* 0.9026* ( cos 44.1)^2

I2 = 0.23274 W/m^2

=======

Comment before rate in case any doubt, will reply for sure.. goodluck